OpenStudy (anonymous):

how do you subtract mixed numbers with a fraction by changing a mixed number to an improper fraction so you are subtracting the improper fraction?

OpenStudy (anonymous):

do you have an example?

OpenStudy (anonymous):

yes 1 1/3 - 1/3 you have to change the mixed number to an improper fraction and subtract by the other fraction

OpenStudy (anonymous):

ok so to make the mixed number into an improper fraction you multiply the denominator by the 1 so 3(1)=3 then you add the numerator 3+1=4 1 1/3=4/3 now you can simply subtract 4/3-1/3=3/3 or 1

OpenStudy (anonymous):

make sense?

OpenStudy (anonymous):

Yeah makes sense

OpenStudy (anonymous):

cool beans:)

OpenStudy (anonymous):

sorry my old account went weird so i had to make one under sammy11 instead of sammo11 ok

OpenStudy (anonymous):

oh that's fine just let me know if you need any more help

OpenStudy (anonymous):

thanks